# Value-Guided Search for Efficient Chain-of-Thought Reasoning

### v1

@@ 3.2 Test-Time Compute Scaling for Search
Figure 3: Test-Time Compute with DeepSeek-VM-1.5B. (Left) Compares best-of-$N$ (BoN), weighted majority voting (WMV) and VGS with either BoN or WMV for the final aggregation. (Right) Compares VGS to majority voting (MV), a standard baseline that does not require a scorer.
@@
Response-Level Selection vs Search-Based Block-Level Selection. While BoN and WMV represent standard approaches for selecting responses using an ORM, block-wise VGS guides response generation through sequential block-by-block selection. As Fig. 3 (left) illustrates, WMV consistently outperforms BoN across all inference budget scales, which demonstrates the benefits of combining MV with value scores. Furthermore, VGS (with WMV as a final aggregation step) yields additional improvements beyond WMV alone. This confirms the benefits of search and aligns with conclusions from previous studies [6, 36, 25]. Interestingly, we do not observe the same benefits of search if BoN is used as a final aggregation step, suggesting that WMV is a critical component to VGS.
@@
VGS vs Majority Voting. As Fig. 3 (right) demonstrates, VGS consistently achieves higher accuracy than MV, attaining equivalent performance while requiring substantially lower inference budgets (also shown in Fig. 1 right). Fully closing the gap with the oracle Pass@N curve may require a larger value model trained on more extensive datasets.
DeepSeek-VM-1.5B vs Baseline Scoring Models. Fig. 4 benchmarks DeepSeek-VM-1.5B against existing PRMs and our BT model. We observe that DeepSeek-VM-1.5B consistently delivers superior performance when employed both as an ORM for WMV (left) and as a guidance mechanism for block-wise search (right). Note that we find our BT model to be surprisingly effective as a search guidance model which suggests the importance of our token-level dataset playing an important role in successful downstream search.
@@ 3.3 Scaling up the Generator Model Sizes
In Fig. 5, we scale up our experiments to guide larger 7B and 14B DeepSeek models. Here, we run VGS with the same search parameters using the same value model DeepSeek-VM-1.5B. Although the 7B and 14B DeepSeek models are in theory OOD for our value model [52], which was trained on DeepSeek-1.5B rollouts, we observe that VGS continues to scale without plateauing as test-time compute increases. This provides some evidence that a value model trained with a weaker verifier policy can generalize effectively and guide the CoTs of stronger models. Such generalization is particularly valuable, as it is significantly cheaper to collect training data from smaller $\pi^{\text{ref}}$ models. This form of “weak-to-strong” generalization [10] appears to be a promising direction for future research.
Figure 6: Ablation: Search Block Size. AIME-24 accuracies for beam search (width $2$) with varying block sizes from 16 to 16384. We found 4096 to be optimal across test-time budgets and benchmarks.
@@ 4 Ablation Studies
To investigate the role of key hyperparameters in search, we perform sensitivity analyses of block size and beam width on AIME-24 across varying inference budgets. We also ablate the amount of DVTS parallelism. These tests suggest that there is a consistent choice of search hyperparameters that work well across inference budgets.
@@ 4.1 Different Search Parameters and Methods
Block Size. We perform beam search with width $2$ using search block sizes from 16 to 16384. Fig. 6 shows AIME-24 accuracies across three inference budgets $N$, revealing that the optimal choice of $4096$ stays consistent across different $N$. We see a decline in performance when searching with more fine-grained blocks.
Beam Width. We perform beam search with block size $4096$ using varying beam widths, with breadth-first-search (BFS) being a special case where beam width is equal to $N$. Fig. 7 (left) shows AIME-24 accuracies across five inference budgets, demonstrating that beam width $2$ is consistently optimal across different $N$. We note our optimal beam width is different from prior works’ which found $4$ to work best [6, 36, 25].
DVTS Parallelism. Fig. 7 (right) shows the role of ablating DVTS from VGS. For each inference budget, we report average accuracies without DVTS and with the best DVTS parallelism $M$. We observe that DVTS becomes more effective at higher budgets and scales better than a single search tree, which is consistent with findings from prior works [6]. However, we find that DVTS is never worse than a single search tree even at smaller inference budgets, which is the opposite conclusion reached by prior works [6]. This discrepancy may be explained by the fact that we use WMV to combine the DVTS responses, which seems to be a more robust way to perform DVTS than BoN (used in prior works) given our findings from Fig. 3.
Figure 7: Ablations: Beam-Width and DVTS. (Left) AIME-24 accuracies for beam search with various widths (Section 2.3) across inference budgets $N$. BFS is equivalent to setting width as $N$. We find that the optimal beam width is robust across multiple TTC budgets. (Right) Averaged accuracy for beam $2$ with and without DVTS. For DVTS, we report the best result with parallelism $M>1$ per inference budget $N$, which we find scales better at higher budgets.
@@
Finally, we directly ablate the role of our value model’s guidance during the search process. We perform VGS (w/ same width, block size and DVTS) but randomly select blocks instead of selecting blocks with the highest value. We still aggregate the final beams via WMV with our value model, so the only change is how intermediate blocks are chosen. We call this process “random search”. Thus, if our value model is helpful for search, we should expect VGS to outperform random search. Indeed, Fig. 8 validates this hypothesis. We also evaluate a hybrid approach where half of DVTS’s parallel trees use random search and the other half use VGS. We find that this hybrid approach lands roughly between pure VGS and pure random search, again validating that block-selection from our value model improves over random selection.
@@ 5 Conclusion
In this paper, we introduced block-wise Value-Guided Search (VGS), a simple yet effective strategy for steering long-context CoT reasoning models. We proposed a scalable token-level value model training pipeline that does not require a pre-defined notion of “step” or expensive per-step annotations. We collect a large dataset of reasoning CoTs (OpenR1-VM) and train a lean 1.5B value model (DeepSeek-VM-1.5B), which we show can effectively guide the CoTs of DeepSeek models up to 14B in size. With extensive experiments, we demonstrate that VGS with DeepSeek-VM-1.5B enjoys better test-time compute scaling than standard methods (e.g., majority voting, best-of-$N$) and other scoring models (e.g., existing PRMs and a BT model), achieving a higher performance ceiling while reducing the FLOPs needed to extract the same performance as baseline methods (Fig. 1). Our results point to VGS as a promising approach to scale TTC of emerging reasoning models.
Discussion of Limitations. Our value model is trained exclusively on completions / roll-outs from a lean reasoning model $\pi^{\text{ref}}$ (e.g., DeepSeek-R1-Distill-Qwen-1.5B). As frontier LLMs continue to advance, the distribution of their generated responses may increasingly diverge from our training distribution, potentially degrading scoring and search performance. To maintain optimal performance, new value models will need to be retrained on rollouts from updated generator policies. However, we do not foresee this as a major practical concern given the simplicity and scalability of our pipeline. To facilitate retraining and adaptation to similar verifiable domains, we open-source our codebase and provide a step-by-step recipe in Appendix B for data collection, training and search inference.
@@ Acknowledgment
JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). ZG is supported by LinkedIn-Cornell Grant. Wen Sun is supported by NSF IIS-2154711, NSF CAREER 2339395 and DARPA LANCER: LeArning Network CybERagents. This research is also supported by grants from the National Science Foundation NSF (IIS-1846210, IIS-2107161, and IIS-1724282, HDR-2118310), the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875), DARPA, arXiv, LinkedIn, Google, and the New York Presbyterian Hospital.
@@
Note: In the appendix, we also provide additional empirical results in Appendix C. Two new results are worth highlighting here. First, in Section C.6, we provide test-time scaling results for guiding DeepSeek-1.5B further trained with PPO on our math dataset. We find that VGS improves test-time scaling compared to MV and WMV, which shows that our method nicely complements policy-based RL training. Moreover, in Section C.7, we include three qualitative examples of contrastive blocks that were selected or rejected by our value model during beam search process. We see that our value model prefers blocks with more straightforward logical deductions, yielding more efficient and effective CoT for reasoning.
@@ Appendix A Related Works
Test-time compute (TTC) broadly refers to algorithms that improve problem-solving performance when given more compute (i.e., FLOPs) at test-time. Fig. 9 summarizes the taxonomy of TTC methods. The simplest TTC methods are score-free in the sense that they do not require access to an external scoring model. A notable example is majority voting (MV), which selects the most frequent answer among $N$ responses, breaking ties randomly [14, 46, 9]. Also known as self-consistency, MV can be applied to tasks where the output space is equipped with an equivalence relation, e.g., mathematical formulae that can be symbolically checked for equality. Other score-free TTC methods include sequentially revising the response via CoT prompting [38, 29] and hybrid methods [41].
There are also score-based TTC methods that employ an external scorer. The coarsest type of a scorer is an outcome reward model (ORM), which takes the full prompt and response as input and produces a scalar that measures the quality / correctness of the response. Popular examples of ORMs include Bradley-Terry reward models [8] or LLM-as-a-Judge [4]. ORMs can be used for best-of-$N$ (BoN), which selects the response with the highest score [14, 4]. ORMs can also be used for weighted majority voting (WMV), which generalizes MV where the strength of a response’s vote is proportional to its ORM score. Weighted MV (WMV) typically provides an improvement over vanilla (unweighted) MV [6, 25], which is also what we observe in our experiments (Fig. 3).
@@
These limitations make it difficult to scale PRMs to long-context reasoning models [18], and all these prior works could only evaluate on short-context models with easier benchmarks such as GSM8k [14] and MATH [23]. In contrast, our paper focuses on scaling process-level guidance to long-context reasoning models, and we propose a block-level search method that mitigates the above limitations. We train a token-level value model by collecting rollouts from random solution prefixes, which requires neither a pre-defined notion of step nor per-step annotations. We use our value model to guide a block-wise search process, where the block size is a hyperparameter and we find there exists a consistent choice that works well across inference budgets (Fig. 6). Crucially, we are able to scale our value-guided search (VGS) to long-context DeepSeek models and demonstrate impressive performance and efficiency gains on challenging math competitions (Fig. 1).
@@
We remark that some prior works proposed to use token-level value models to reweight the next-token distribution of the generator [28, 49, 19, 52]. However, these methods require one classifier call per token, which is more expensive than block-wise search. Moreover, token-level guidance might also be less effective because the imperfect value model may introduce cascading errors if queried at every token. We highlight that Mudgal et al. [28] also experimented with block-wise BFS and found this to be more effective at scaling test-time compute than reweighting the next-token distribution (i.e., token-wise guidance). One drawback of block-wise BFS is that the blocks may all become correlated due to sharing the same prefix. Thus, we build upon Mudgal et al. [28] by proposing to use beam search, which we show yields better test-time scaling for reasoning models (Fig. 6).
@@ Appendix B Summary of VGS Pipeline
    Train a classifier / value model by following (Section 2.1). Sweep hyperparameters such as learning rate.
@@ C.1 Full Main Results Table
We reproduce Table 1 with additional baselines: DeepSeek-7B, DeepSeek-14B, o1-mini-medium.
@@
C.2 Per-benchmark Plots for Fig. 3
Figure 10: Per-benchmark results for Fig. 3. (Left) Compares best-of-$N$ (BoN), weighted majority voting (WMV) and VGS with either BoN or WMV for the final aggregation. (Right) Compares VGS to majority voting (MV), a standard baseline that does not require a scorer.
C.3 Per-benchmark Plots for Fig. 4
Figure 11: Per-benchmark results for Fig. 4. Comparison of our 1.5B value model (VM), our 1.5B Bradley-Terry reward model (BT), and two 7B state-of-the-art PRMs for two TTC scaling methods: (Left) WMV or (Right) VGS (with WMV as a final aggregation step).
C.4 Per-benchmark Plots for Fig. 5
Figure 12: Per-benchmark results for Fig. 5. With the same DeepSeek-VM-1.5B providing guidance, search continues to improve with more test-time compute.
C.5 Per-benchmark Plots for Fig. 8
Figure 13: Per-benchmark results for Fig. 8. Random search is the same search process as VGS except intermediate blocks are randomly selected instead of using our value model.
@@ C.6 Results for Guiding a PPO Policy
Table 3 presents the comparison between the DeepSeek-1.5B model and the PPO-trained model (DeepSeek-1.5B-PPO). As N increases, the performance gap gradually narrows. While the PPO-trained model performs competitively at lower $N$ values, it is surpassed by the base model at Pass@32 on both the AIME-24 and HMMT-25 datasets. This decline in performance could be attributed to the reduced entropy of the model after PPO training, which limits the diversity of model generations and negatively impacts performance at higher Pass@N.
We report our value-guided results of the PPO-trained model in Fig. 14. We observe that VGS nicely complements PPO training and provides additional test-time compute gains in performance compared to WMV and MV.

### v2

@@ 3.2 Test-Time Compute Scaling for Search
Figure 3: Test-Time Compute with DeepSeek-VM-1.5B. (Left) Compares best-of-NN (BoN), weighted majority voting (WMV) and VGS with either BoN or WMV for the final aggregation. (Right) Compares VGS to majority voting (MV), a standard baseline that does not require a scorer.
@@
Response-Level Selection vs Search-Based Block-Level Selection. While BoN and WMV represent standard approaches for selecting responses using an ORM, block-wise VGS guides response generation through sequential block-by-block selection. As Fig.˜3 (left) illustrates, WMV consistently outperforms BoN across all inference budget scales, which demonstrates the benefits of combining MV with value scores. Furthermore, VGS (with WMV as a final aggregation step) yields additional improvements beyond WMV alone. This confirms the benefits of search and aligns with conclusions from previous studies [6, 37, 26]. Interestingly, we do not observe the same benefits of search if BoN is used as a final aggregation step, suggesting that WMV is a critical component to VGS.
@@
VGS vs Majority Voting. As Fig.˜3 (right) demonstrates, VGS consistently achieves higher accuracy than MV, attaining equivalent performance while requiring substantially lower inference budgets (also shown in Fig.˜1 right). Fully closing the gap with the oracle Pass@N curve may require a larger value model trained on more extensive datasets.
DeepSeek-VM-1.5B vs Baseline Scoring Models. Fig.˜4 benchmarks DeepSeek-VM-1.5B against existing PRMs and our BT model. We observe that DeepSeek-VM-1.5B consistently delivers superior performance when employed both as an ORM for WMV (left) and as a guidance mechanism for block-wise search (right). Note that we find our BT model to be surprisingly effective as a search guidance model which suggests the importance of our token-level dataset playing an important role in successful downstream search.
@@ 3.3 Scaling up the Generator Model Sizes
In Fig.˜5, we scale up our experiments to guide larger 7B and 14B DeepSeek models. Here, we run VGS with the same search parameters using the same value model DeepSeek-VM-1.5B. Although the 7B and 14B DeepSeek rollins are OOD for our value model (trained on 1.5B traces), we observe that VGS continues to scale without plateauing as test-time compute increases. This provides some evidence that a value model trained with a weaker verifier policy can generalize effectively and guide the CoTs of stronger models, which is useful since it is much cheaper to collect training data from smaller $\pi^{\text{ref}}$ models. This form of “weak-to-strong” generalization [10] appears to be a promising direction for future research.
Figure 6: Ablation: Search Block Size. AIME-24 accuracies for beam search (width 22) with varying block sizes from 16 to 16384. We found 4096 to be optimal across test-time budgets and benchmarks.
@@ 4 Ablation Studies
To investigate the role of key hyperparameters in value-guided search, we perform ablation analyses of block size and beam width on AIME-24 across varying inference budgets. We also ablate the amount of DVTS parallelism. These tests suggest that there is a consistent choice of search hyperparameters that work well across inference budgets.
@@ 4.1 Different Search Parameters and Methods
Block Size. We perform beam search with width $2$ using search block sizes from 16 to 16384. Fig.˜6 shows AIME-24 accuracies across three inference budgets $N$, revealing that the optimal choice of $4096$ stays consistent across different $N$. We see a decline in performance when searching with more fine-grained blocks.
Beam Width. We perform beam search with block size $4096$ using varying beam widths, with breadth-first-search (BFS) being a special case where beam width is equal to $N$. Fig.˜7 (left) shows AIME-24 accuracies across five inference budgets, demonstrating that beam width $2$ is consistently optimal across different $N$. We note our optimal beam width is different from prior works’ which found $4$ to work best [6, 37, 26].
DVTS Parallelism. Fig.˜7 (right) shows the role of ablating DVTS from VGS. For each inference budget, we report average accuracies without DVTS and with the best DVTS parallelism $M$. We observe that DVTS becomes more effective at higher budgets and scales better than a single search tree, which is consistent with findings from prior works [6]. However, we find that DVTS is never worse than a single search tree even at smaller inference budgets, which is the opposite conclusion reached by prior works [6]. This discrepancy may be explained by the fact that we use WMV to combine the DVTS responses, which seems to be a more robust way to perform DVTS than BoN (used in prior works) given our findings from Fig.˜3.
Figure 7: Ablations: Beam-Width and DVTS. (Left) AIME-24 accuracies for beam search with various widths (Section˜2.3) across inference budgets NN. BFS is equivalent to setting width as NN. We find that the optimal beam width is robust across multiple TTC budgets. (Right) Averaged accuracy for beam 22 with and without DVTS. For DVTS, we report the best result with parallelism $M>1$ per inference budget NN, which we find scales better at higher budgets.
@@
Finally, we directly ablate the role of our value model’s guidance during the search process. We perform VGS (w/ same width, block size and DVTS) but randomly select blocks instead of selecting blocks with the highest value. We still aggregate the final beams via WMV with our value model, so the only change is how intermediate blocks are chosen. We call this process “random search”. Thus, if our value model is helpful for search, we should expect VGS to outperform random search. Indeed, Fig.˜8 validates this hypothesis. We also evaluate a hybrid approach where half of DVTS’s parallel trees use random search and the other half use VGS. We find that this hybrid approach lands roughly between pure VGS and pure random search, again validating that block-selection from our value model improves over random selection.
@@ 5 Conclusion
In this paper, we introduced block-wise Value‐Guided Search (VGS), a simple yet effective strategy for steering long‐context CoT reasoning models. We proposed a scalable token‐level value model training pipeline that does not require a pre-defined notion of “step” or expensive per‐step annotations. We collect a large dataset of reasoning CoTs (OpenR1-VM) and train a lean 1.5B value model (DeepSeek-VM-1.5B), which we show can effectively guide the CoTs of DeepSeek models up to 14B in size. With extensive experiments, we demonstrate that VGS with DeepSeek-VM-1.5B enjoys better test-time compute scaling than standard methods (e.g., majority voting, best-of-$N$) and other scoring models (e.g., existing PRMs and a BT model), achieving a higher performance ceiling while reducing the FLOPs needed to extract the same performance as baseline methods (Fig.˜1). Our results point to VGS as a promising approach to scale TTC of emerging reasoning models.
Discussion of Limitations. Our value model is trained exclusively on completions / roll-outs from a lean reasoning model $\pi^{\text{ref}}$ (e.g., DeepSeek-R1-Distill-Qwen-1.5B). As frontier LLMs continue to advance, the distribution of their generated responses may increasingly diverge from our training distribution, potentially degrading scoring and search performance. To maintain optimal performance, new value models will need to be retrained on rollouts from updated generator policies. However, we do not foresee this as a major practical concern given the simplicity and scalability of our pipeline. To facilitate retraining and adaptation to similar verifiable domains, we open-source our codebase and provide a step-by-step recipe in Appendix˜B for data collection, training and search inference.
@@ Acknowledgment
KW is supported by a Google PhD Fellowship. JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). ZG is supported by LinkedIn-Cornell Grant. Wen Sun is supported by NSF IIS-2154711, NSF CAREER 2339395 and DARPA LANCER: LeArning Network CybERagents. This research is also supported by grants from the National Science Foundation NSF (IIS-1846210, IIS-2107161, and IIS-1724282, HDR-2118310), the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875), DARPA, arXiv, LinkedIn, Google, and the New York Presbyterian Hospital.
@@
Note: In the appendix, we also provide additional empirical results in Appendix˜C. Two new results are worth highlighting here. First, in Section˜C.6, we provide test-time scaling results for guiding DeepSeek-1.5B further trained with PPO on our math dataset. We find that VGS improves test-time scaling compared to MV and WMV, which shows that our method nicely complements policy-based RL training. Moreover, in Section˜C.7, we include three qualitative examples of contrastive blocks that were selected or rejected by our value model during beam search process. We see that our value model prefers blocks with more straightforward logical deductions, yielding more efficient and effective CoT for reasoning.
@@ Appendix A Related Works
Test-time compute (TTC) broadly refers to algorithms that improve problem-solving performance when given more compute (i.e., FLOPs) at test-time. Fig.˜9 summarizes the taxonomy of TTC methods. The simplest TTC methods are score-free in the sense that they do not require access to an external scoring model. A notable example is majority voting (MV), which selects the most frequent answer among $N$ responses, breaking ties randomly [14, 47, 9]. Also known as self-consistency, MV can be applied to tasks where the output space is equipped with an equivalence relation, e.g., mathematical formulae that can be symbolically checked for equality. Other score-free TTC methods include sequentially revising the response via CoT prompting [39, 30] and hybrid methods [42].
There are also score-based TTC methods that employ an external scorer. The coarsest type of a scorer is an outcome reward model (ORM), which takes the full prompt and response as input and produces a scalar that measures the quality / correctness of the response. Popular examples of ORMs include Bradley-Terry reward models [8] or LLM-as-a-Judge [4]. ORMs can be used for best-of-$N$ (BoN), which selects the response with the highest score [14, 4]. ORMs can also be used for weighted majority voting (WMV), which generalizes MV where the strength of a response’s vote is proportional to its ORM score. Weighted MV (WMV) typically provides an improvement over vanilla (unweighted) MV [6, 26], which is also what we observe in our experiments (Fig.˜3).
@@
These limitations make it difficult to scale PRMs to long-context reasoning models [19], and all these prior works could only evaluate on short-context models with easier benchmarks such as GSM8k [14] and MATH [24]. In contrast, our paper focuses on scaling process-level guidance to long-context reasoning models, and we propose a block-level search method that mitigates the above limitations. We train a token-level value model by collecting rollouts from random solution prefixes, which requires neither a pre-defined notion of step nor per-step annotations. We use our value model to guide a block-wise search process, where the block size is a hyperparameter and we find there exists a consistent choice that works well across inference budgets (Fig.˜6). Crucially, we are able to scale our value-guided search (VGS) to long-context DeepSeek models and demonstrate impressive performance and efficiency gains on challenging math competitions (Fig.˜1).
@@
We remark that some prior works proposed to use token-level value models to reweight the next-token distribution of the generator [29, 50, 20, 53]. However, these methods require one classifier call per token, which is more expensive than block-wise search. Moreover, token-level guidance might also be less effective because the imperfect value model may introduce cascading errors if queried at every token. We highlight that Mudgal et al. [29] also experimented with block-wise BFS and found this to be more effective at scaling test-time compute than reweighting the next-token distribution (i.e., token-wise guidance). One drawback of block-wise BFS is that the blocks may all become correlated due to sharing the same prefix. Thus, we build upon Mudgal et al. [29] by proposing to use beam search, which we show yields better test-time scaling for reasoning models (Fig.˜6).
Recently, Fu et al. [18] showed that confidence as measured by the average next-token entropy along the thinking trace can also be used to guide the CoT generation process to improve TTC performance. An advantage of Fu et al. [18] is that it does not require training a value model and can be applied directly given a generator policy. However, it does not leverage any reward signals and may fall short in planning long-term strategies [53]. It is a promising future direction to explore how to combine uncertainty-based methods with value-based methods.
@@ Appendix B Summary of VGS Pipeline
    Train a classifier / value model by following (Section˜2.1). Sweep hyperparameters such as learning rate.
@@ C.1 Full Main Results Table
We reproduce Table˜1 with additional results and baselines. The results with closed source models were obtained via API calls in May 2025. We see that with a budget of $256$, our 1.5B value model can guide DeepSeek-1.5B (total parameter count is 3B) to reach parity with the pass@1 of o3-mini-medium.
@@
C.2 Per-benchmark Plots for Fig.˜3
Figure 10: Per-benchmark results for Fig.˜3. (Left) Compares best-of-NN (BoN), weighted majority voting (WMV) and VGS with either BoN or WMV for the final aggregation. (Right) Compares VGS to majority voting (MV), a standard baseline that does not require a scorer.
C.3 Per-benchmark Plots for Fig.˜4
Figure 11: Per-benchmark results for Fig.˜4. Comparison of our 1.5B value model (VM), our 1.5B Bradley-Terry reward model (BT), and two 7B state-of-the-art PRMs for two TTC scaling methods: (Left) WMV or (Right) VGS (with WMV as a final aggregation step).
C.4 Per-benchmark Plots for Fig.˜5
Figure 12: Per-benchmark results for Fig.˜5. With the same DeepSeek-VM-1.5B providing guidance, search continues to improve with more test-time compute.
C.5 Per-benchmark Plots for Fig.˜8
Figure 13: Per-benchmark results for Fig.˜8. Random search is the same search process as VGS except intermediate blocks are randomly selected instead of using our value model.
@@ C.6 Results for Guiding a PPO Policy
Table˜3 presents the comparison between the DeepSeek-1.5B model and the PPO-trained model (DeepSeek-1.5B-PPO). As N increases, the performance gap gradually narrows. While the PPO-trained model performs competitively at lower $N$ values, it is surpassed by the base model at Pass@32 on both the AIME-24 and HMMT-25 datasets. This decline in performance could be attributed to the reduced entropy of the model after PPO training, which limits the diversity of model generations and negatively impacts performance at higher Pass@N.
We report our value-guided results of the PPO-trained model in Fig.˜14. We observe that VGS nicely complements PPO training and provides additional test-time compute gains in performance compared to WMV and MV.